OpenStudy (anonymous):

Find the LCD for the following fractions:

OpenStudy (anonymous):

\[\frac{ 3a^2 }{ a^2+a+1 } , \frac{ 2a }{ a-1 } , \frac{ a^3 }{ a^3-1 }\]

OpenStudy (anonymous):

Well, you can always find "a" common denominator by multiplying all the denominator's together. I haven't convinced myself that this is the Least Common Denominator in this case, however.

OpenStudy (anonymous):

For instance: \[\frac{3a^2}{a^2+a+1} = \frac{(3a^2)(a-1)(a^3-1)}{(a^2+a+1)(a-1)(a^3-1)}\]

OpenStudy (anonymous):

I don't understand..lol

OpenStudy (anonymous):

Well, we can always multiply a number by 1, because this doesn't change the value. So, if we multiply by the same thing on the top and bottom of a fraction, that's the same as multiplying by 1. So we multiply each fraction on the top and bottom by the denominator of each of the other fractions (as I have shown above)

OpenStudy (anonymous):

If we do this to each of the fractions, they'll all have the same denominator.